myininaya (myininaya):

Here is a problem that wolfram won't evaluate that someone had a question about. \[int_{}^{} sqrt{1+\frac{1}{4} \cdot frac{x^2 (x-2)^2}{(1-x)^2}} dx\]

myininaya (myininaya):

\[\int\limits_{}^{} \sqrt{1+\frac{1}{4} \cdot \frac{x^2 (x-2)^2}{(1-x)^2}} dx \]

myininaya (myininaya):

I honestly don't see a way by elementary ways.

OpenStudy (anonymous):

JamesJ can do it!

hero (hero):

Let's see if my calc evaluate it :P

OpenStudy (anonymous):

Otherwise I will try sympy

OpenStudy (cwrw238):

breaking it into partial fractions is a possibility

hero (hero):

It doesn't come out to anything simple at all

OpenStudy (earthcitizen):

partial fractions then substitution

OpenStudy (jamesj):

Ok, the integrand is equal to \[ \frac{1}{2} \sqrt{ \frac{ 4(1-x)^2 + x^2(x-2)^2 }{ (1-x)^2 } } = \frac{1}{2} \sqrt{ \frac{ (x^2 - 2x + 2)^2}{(1-x)^2} } \] \[ = \frac{1}{2} \frac{ x^2 - 2x + 2}{1-x} = -x - \frac{1}{x-1} + 1\]

hero (hero):

Well, someone said james could do it and he did it

myininaya (myininaya):

why didn't i see that?

OpenStudy (jamesj):

I'm just happy to find another example of Wolfram Alpha defeated by pencil and paper. :-)

myininaya (myininaya):

good job another point for the human team

OpenStudy (anonymous):

pencil and paper > modern CAS

hero (hero):

TI-Nspire CAS defeated too

OpenStudy (mr.math):

Apparently James is traditionalist! ;D

OpenStudy (anonymous):

I couldn't figure out how to use sympy :-P

OpenStudy (jamesj):

**Correction: I dropped the half in the last expression.

myininaya (myininaya):

that's cool who cares about the constant multiple we put it in a form easy to integrate well you did anyways

OpenStudy (jamesj):

The factorization of the numerator isn't obvious. Don't feel bad. But I do think this problem is a bit silly. What was the motivation for that integrand in that form? I'd bet a nickel the integrand was constructed from the simplified form as a puzzle, rather than being motivated by any other problem in mathematics or the sciences.

myininaya (myininaya):

lol i made a puzzle by accident

myininaya (myininaya):

the original problem was find the length of \[y=\frac{x}{2}-\frac{x^2}{4}+\frac{1}{2}\ln(1-x) \text{ on interval } [0, .5]\] so i found y' which is \[y'=\frac{1}{2}(1-x-\frac{1}{x-1})\] then i wrote (y')^2 as \[(y')^2=\frac{1}{4} \cdot \frac{x^2(x-2)^2}{(1-x)^2}\] which made it uglier than it should had been

myininaya (myininaya):

wasn't thinking

OpenStudy (jamesj):

No, I owe you a nickel. This did come from a real problem.

myininaya (myininaya):

What do you mean?

OpenStudy (jamesj):

I mean the integral came from a real problem, vs. just being a slightly annoying puzzle.

myininaya (myininaya):

Ok. lol

myininaya (myininaya):

so what was your trick for factoring the numerator?

OpenStudy (anonymous):

he used... wolfram!

myininaya (myininaya):

no james doesn't use wolfram he doesn't believe in it lol

OpenStudy (anonymous):

right

OpenStudy (jamesj):

I knew I wanted a perfect square to get rid of the square trot. So I went looking for it. As ever with three things squared, I think it's easier to find it as \[ ((a+b) + c)^2 \] vs as \[ (a+b+c)^2 \] So after I realized it wasn't (a+b)^2, I went looking for things in that form.

OpenStudy (jamesj):

*square root

OpenStudy (asnaseer):

JamesJ - this is the 2nd such problem that wolfram couldn't deal with. should these be reported to the wolfram team so that they can improve their algorithms?

OpenStudy (jamesj):

Geez asnaneer. Just send me to the glue factory already. ;-)

OpenStudy (anonymous):

the integral algorithm is probably complicated :-D

myininaya (myininaya):

James thanks you are the coolest i think i learned a little

OpenStudy (asnaseer):

:-) - just trying to improve the mathematical community - thats all...

OpenStudy (anonymous):

I bet that on JamesJ's copy of the monograph on the Riemann Hypothesis is a note on the margin saying : "I have discovered a truly marvellous proof of this, which this margin is too narrow to contain."

OpenStudy (asnaseer):

BTW JamesJ - what type of glue do you prefer? :-D

OpenStudy (jamesj):

I'm sure Wolfram could find it, but it has to make decisions about how long to run certain algorithms. With a tweak of its "time out" parameters (or @agd, tree pruning heuristic), it would find this solution.

OpenStudy (anonymous):

the wolfram solution is not as elegant as JamesJ's solution though. I think it's an interesting exercise to do a little research on how those symbolic integration/factoring stuff work

OpenStudy (jamesj):

I dread to think how it found that solution, because it suggests quite a tortured process. But the real problem is once it did, it stopped and it didn't try and simplify the answer, which clearly it could have. But in any case, it looks like it didn't use the procedure I did: try and find a perfect square and rid of the square root asap. Anyway, time to celebrate by making a cup of tea. ;-)

OpenStudy (asnaseer):

I agree - plus the wolfram solution doesn't really improve your understanding of the problem.

myininaya (myininaya):

lol

OpenStudy (anonymous):

let's see what that function returns when I try to evaluate myininaya's integral

OpenStudy (anonymous):

still running :-P

OpenStudy (anonymous):

http://ideone.com/9Cxc3 here's the integral method again

OpenStudy (anonymous):

I think we should check in 50 minutes for the output of the integral method. :-D

OpenStudy (anonymous):

\[# ______________________ / 2 2 / 0.25*x *(x - 2) / ---------------- + 1 / 2 \/ (-x + 1) \] was what I gave it,

OpenStudy (anonymous):

It couldn't do it. It returned :\ / | | _________________ | / 2 2 | / x *(x - 2) | / ----------- + 1 dx | / 2 | \/ 4*(-x + 1) | /